# Cross-talk in superconducting qubit lattices with tunable couplers – comparing transmon and fluxonium architectures

### v1

@@ Abstract
Cross-talk between qubits is one of the main challenges for scaling superconducting quantum processors. Here, we use the density-matrix renormalization-group to numerically analyze lattices of superconducting qubits from a perspective of many-body localization. Specifically, we compare different architectures that include tunable couplers designed to decouple qubits in the idle state, and calculate the residual $ZZ$ interactions as well as the inverse participation ratio in the computational basis states. For transmon qubits outside of the straddling regime, the results confirm that tunable C-shunt flux couplers are significantly more efficient in mitigating the $ZZ$ interactions than tunable transmons. A recently proposed fluxonium architecture with tunable transmon couplers is demonstrated to also maintain its strong suppression of the $ZZ$ interactions in larger systems, while having a higher inverse participation ratio in the computational basis states than lattices of transmon qubits. Our results thus suggest that fluxonium architectures may feature lower cross talk than transmon lattices when designed to achieve similar gate speeds and fidelities.
@@ I Introduction
The rest of the paper is organized as follows: Section II introduces the model for the superconducting qubit systems and the theoretical approach. In Sec. III, we analyze transmon systems coupled in three different ways: directly via capacitive coupling, via an additional tunable transmon, and using a C-shunt flux coupler, while Sec. IV deals with fluxonium systems coupled via Transmon tunable couplers. The results are summarized in Sec. V.
@@ II.1 System Hamiltonian
where $\hat{H}_{\text{loc}}$ is the sum of the local terms, $T$ and $T_{c}$ are the coupling strengths between the qubits and between the qubits and the coupler, respectively, $\hat{n}_{i}$ is the Cooper-pair number operator for the qubit or coupler with index $i$, and $\langle q_{1}q_{2}c\rangle$ denotes a pair of qubits and the corresponding coupler $c$. Equation (1) should be regarded as an approximation of a circuit with coupling capacitances between the qubits and couplers connected by lines in Fig. 1. A more accurate treatment would involve extracting the couplings from the inverse of the capacitive matrix, including longer-ranged terms that we neglect here, since they are typically at least an order of magnitude smaller than the couplings considered in Eq. (1).
Depending on its architecture, the tunable coupler can couple to its neighboring qubits either with the same sign $(s=1)$ or with opposite signs $(s=-1)$. While floating architectures with $s=-1$ are preferred for C-shunt flux couplers (see  [12]), we also analyze floating transmon coupler architectures that couple to qubits with opposite signs. This allows for a more direct comparison between both architectures. For the fluxonium system, we adopt the architecture described in  [14], where the transmon couplers couple to fluxonium qubits with the same sign. The local terms $\hat{H}_{\text{loc}}$ depend on the type of qubit or coupler and are discussed in the corresponding sections. They can also include disorder, which leads to many-body localization (MBL) of the system [17].
@@ II.2 MBL states in superconducting architectures
$\hat{H}_{\text{MBL}}=\sum_{n\in\{0,1\}^{N}}w_{n}\bigotimes_{i=1}^{N}\hat{\tau}_{i}^{n_{i}}\,,$ (2)
where $N$ is the number of sites, $w_{n}$ are real coefficients and $l$-bits $\hat{\tau}_{i}$ are operators related to the Pauli $Z$ operators by a quasi-local unitary transformation. Bitstrings $n$ of length $N$ are used to label the different terms in the operator expansion, with $n_{i}=1$ indicating a $\hat{\tau}$ operator at site $i$. The definition of the $l$-bit operators is not unique. One way to construct a set of $l$-bits is to diagonalize the Hamiltonian, define an injective map from the product states $|n\rangle=\otimes_{i}|n_{i}\rangle$ to the eigenstates $|\tilde{n}\rangle$ that maximizes the sum of the overlaps $\sum_{n\in\{0,1\}^{N}}|\langle n|\tilde{n}\rangle|^{2}$ [21], and set $\hat{\tau}_{i}=\sum_{n}(-1)^{n_{i}}|\tilde{n}\rangle\langle\tilde{n}|$. Alternatively, one may use a mapping based on flow equations [22] or tracing the energy levels starting from a decoupled limit with maximum localization [7].
@@
Models of superconducting qubits like Eq. (1) also fit into the above framework, with $|\tilde{n}\rangle$ becoming the computational basis states and $\hat{\tau}_{i}$ the $Z$ operators for the dressed qubits. A difference compared with the spin-1/2 case is that the systems include higher-lying excited states for the qubits, and tunable couplers as additional sites. The computational basis states are therefore only those eigenstates that are matched with the product states in which the qubits are in one of the two lowest energy levels and the couplers are in their ground states (ignoring coupling terms between qubits or couplers). After projecting onto the $2^{N}$-dimensional computational subspace, the Hamiltonian again has the form (2), which defines the longitudinal interactions between idling qubits.
@@
In addition to these interactions, we want to investigate how well the computational basis is localized, i.e., how much the dressed basis states $|\tilde{n}\rangle$ differ from the corresponding product states $|n\rangle$. A possible measure for this is the inverse participatio ratio (IPR). For a state $|\psi\rangle$, it is defined as
IPR $=\sum_{n}|\langle\psi|n\rangle|^{4}\,,$ (3)
where the sum is over all states in the undressed basis.
@@
To determine the coupling coefficients $w_{n}$ and the IPR numerically, we make use of the fact that MBL states in one dimension admit an efficient approximation as MPS [23] and employ the DMRG-X method [16]. The difference compared to the regular DMRG method is that it targets eigenstates based on their overlap with a reference state, in this case the relevant product states $|n\rangle$. Unless noted otherwise, the mapping between product states and eigenstates simplifies to picking the eigenstate with the largest overlap, and one DMRG-X calculation for each $n$ is sufficient. If two product states map to the same eigenstate, we do additional DMRG-X simulations with an orthogonalization against the previously calculated states, and switch to the more general mapping mentioned above that maximizes the sum of overlaps [21].
@@
The interaction coefficients are obtained from the energies $E_{n}$ by a Walsh-Hadamard transformation [7]
$w_{n}$ $=\frac{1}{2^{N}}\sum_{n^{\prime}\in\{0,1\}^{N}}(-1)^{\sum_{i}n_{i}n_{i}^{\prime}}E_{n^{\prime}}\,.$ (4)
@@
Because of the larger local Hilbert spaces, we used a single-site variant of the DMRG-X algorithm. As long as the bond dimension $\chi$ is small, the local eigenproblems in the DMRG-X can be efficiently solved by applying a standard solver to the dense-matrix representation of the projected Hamiltonians. For larger $\chi$, we switched to a matrix-free Krylov method and look for the eigenvector to the smallest eigenvalue of the $(\hat{H}-\lambda)^{2}$, where $\lambda$ is the approximate energy from a calculation with small $\chi$. We checked the convergence for each state by monitoring the variance $\langle\tilde{n}|\hat{H}^{2}|\tilde{n}\rangle-\langle\tilde{n}|\hat{H}|\tilde{n}\rangle^{2}$. In cases where the calculations did not converge to the targeted eigenstate with sufficient accuracy straightaway, usually because the DMRG simulations got stuck in a superposition of two eigenstates close in energy, we found it helpful to apply an iteration of the shift-invert DMRG method [24], implemented with the library Krylov.jl [25]. To expand the MPS bonds, we used the method described in Ref. [26].
@@ III Transmon qubits
$\hat{H}_{\text{T}}$ $=4E_{C}\hat{n}^{2}-E_{J}\cos(\hat{\phi})\,,$ (5)
where $\hat{\phi}$ is the superconducting phase that obeys the commutation relation $[\hat{\phi},\hat{n}]=\text{i}$ with the Cooper-pair number operator $\hat{n}$. It is convenient to also introduce the ladder operators for the quadratic part of $\hat{H}_{\text{T}}$ via
@@
$\hat{H}_{\text{T}}$ $\approx\omega\,\hat{b}^{\dagger}\hat{b}+\frac{\alpha}{2}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b}$ (8)
with $\omega=\sqrt{8E_{C}E_{J}}-E_{C}$ and $\alpha=-E_{C}$. In the following we truncate the Hilbert space to the lowest $6$ eigenstates and study lattices of superconducting qubits described by the Hamiltonian (1) with local terms as in Eq. (8). The charging energy is assumed to be $E_{C}=0.2\,\text{GHz}$ for all qubits, while the Josephson energy $E_{J}$ is site-dependent, introducing disorder in the qubit frequencies $\omega$. Throughout this work, we set $\hbar=1$
@@
$\omega_{i}$ $=\bar{\omega}+\Delta\sqrt{2}\sin\Big{[}\pi x_{i}\big{(}y_{i}+\sqrt{y_{i}^{2}+4}\big{)}\Big{]}\,.$ (9)
In an infinite ladder, $\bar{\omega}$ and $\Delta$ become the average frequency and the standard deviation, respectively. We choose $\bar{\omega}=5.8\,\text{GHz}$, $\Delta=0.33\,\text{GHz}$, and an $8$-qubit lattice with $x_{i}\in\{1,2,3,4\}$ and $y_{i}\in\{1,2\}$ (see Table 1). For the couplings between transmon qubits, we consider three different cases: the ladder (i) with only static capacitive coupling, (ii) with tunable transmons and (iii) with tunable C-shunt flux couplers.
qubit $E_{C}$ $E_{J}$ $\omega$
$1$ 0.2 24.33 6.040
$2$ 0.2 19.52 5.389
$3$ 0.2 26.13 6.266
$4$ 0.2 19.68 5.411
$5$ 0.2 20.20 5.485
$6$ 0.2 26.12 6.265
$7$ 0.2 19.81 5.430
$8$ 0.2 23.11 5.881
@@ III.1 Static capacitive coupling
[…]
Figure 2: (a) Interaction coefficients $w_{n}$ for transmons without couplers on a ladder. The graphical labels indicate the positions of the involved Z operators (up to rotations and translations). In the shaded region, the state assignment is ambiguous. The dips in some curves signal sign changes of the corresponding interaction terms. (b) Average IPR for all computational basis states and separated based on the states of the qubits $3$ and $6$.
It is expected that resonant clusters of sites form as an MBL system gets closer to the delocalization transition [31]. Here, the first sign of this appears in interaction coefficient for the next-nearest neighbor qubits 3 and 6, whose frequencies $\omega_{i}$ are almost degenerate. An indication of the strong hybridization is that the IPR is much lower for states with a single excitation between the two qubits. We also observe substantial long-ranged couplings $w_{n}$ involving qubits 3 and 6 that correspond to the bitstrings $n\in\{00101000,00010100,00111000,00011100\}$ in Fig. 2(a). For $T>15$ MHz, the definition of the computational basis states becomes ambiguous, as picking the eigenstate with the largest overlap is no longer an injective map. The results shown are based on the state assigment described in Sec. II. There is still an overall increase of the interaction strengths $|w_{n}|$ with $T$, but with jumps for some coefficients due to changes in the state mapping. One could also assign the computational basis states by tracing the spectrum as a function of $T$ and identifying the anticrossings where the state labels switch [7]. Since that method is based on the energy, it might lead to a less erratic $T$-dependence of the interaction coefficients for large capacitive couplings. It is clear, however, that the system in this regime would not be suitable as a quantum-computing device and that any method of state assignment breaks down for sufficiently large $T$ when the system becomes delocalized.
@@ III.2 Tunable transmon couplers
As shown in Fig. 3, the strength of the nearest-neighbor terms is reduced by about an order of magnitude compared to the system without tunable couplers, and there is a clearer separation in energy scales to longer-ranged interactions. Specific long-ranged terms related to the qubits 3 and 6 still become significant for large enough $T$, however, indicating that these tunable couplers are not able to fully remove the harmful effect of nearly resonant qubits. At very strong couplings $T=20\,$MHz, there are jumps in some of the interaction coefficients $w_{n}$, likely caused by the switching of state labels at avoided level crossings [7]. Such difficulties in assigning the computational basis states are pushed to larger $T$ compared to the system without tunable couplers.
@@
with $S(\rho)=-\text{Tr}(\rho\ln\rho)$ the von Neumann entanglement entropy and $\rho_{X}=\text{Tr}_{\bar{X}}(|\psi\rangle\langle\psi|)$ the reduced density matrix of subsystem $X$ (with complement $\bar{X}$). This quantity, averaged over all computational basis states and qubits, does not increase much until around $T=10\,$MHz, whereas the corresponding quantity in the model without couplers already starts to rise around $T=5\,$MHz. Tunable couplers thus appear to be effective in keeping the qubits localized in the sense that they reduce the spread of the actual (slighly hybridized) computational basis to other qubit sites.
Figure 4: (a) IPR averaged over all states of the computational basis for a ladder with either tunable transmons or C-shunt flux qubits as couplers. The inset shows the average hopping amplitudes between qubits and couplers, taking the factors in Eqs. (7) and  (19) into account. (b) averaged mutual information between a qubit and the remaining system excluding the connected couplers.

### v2

@@ Abstract
Cross-talk between qubits is one of the main challenges for scaling superconducting quantum processors. Here, we use the density-matrix renormalization group to numerically analyze lattices of superconducting qubits from a perspective of many-body localization. Specifically, we compare different architectures that include tunable couplers designed to decouple qubits in the idle state, and calculate the residual $ZZ$ interactions as well as the inverse participation ratio in the computational basis states. For transmon qubits outside of the straddling regime, the results confirm that tunable C-shunt flux couplers are significantly more efficient in mitigating the $ZZ$ interactions than tunable transmons. A recently proposed fluxonium architecture with tunable transmon couplers is demonstrated to also maintain its strong suppression of the $ZZ$ interactions in larger systems, while having a higher inverse participation ratio in the computational basis states than lattices of transmon qubits. Our results thus suggest that fluxonium architectures may feature lower cross talk than transmon lattices when designed to achieve similar gate speeds and fidelities.
@@ I Introduction
The rest of the paper is organized as follows: Section II introduces the model for the superconducting qubit systems and the theoretical approach. In Sec. III, we analyze transmon systems coupled in three different ways: directly via capacitive coupling, via an additional tunable transmon, and using a C-shunt flux coupler, while Sec. IV deals with fluxonium systems coupled via transmon tunable couplers. We end with a discussion of the results and a summary in Secs. V and VI, respectively.
@@ II.1 System Hamiltonian
where $\hat{H}_{\text{loc}}$ is the sum of the local terms, $T$ and $T_{c}$ are the coupling strengths between the qubits and between the qubits and the coupler, respectively, $\hat{n}_{i}$ is the Cooper-pair number operator for the qubit or coupler with index $i$, and $\langle q_{1}q_{2}c\rangle$ denotes a pair of qubits and the corresponding coupler $c$. Equation (1) should be regarded as an approximation of a circuit with coupling capacitances between the qubits and couplers connected by lines in Fig. 1. A more accurate treatment would involve extracting the couplings from the inverse of the capacitance matrix, including longer-ranged terms that we neglect here, since they are typically at least an order of magnitude smaller than the couplings considered in Eq. (1).
Depending on its architecture, the tunable coupler can couple to its neighboring qubits either with the same sign $(s=1)$ or with opposite signs (s=1)(s=-1). While floating architectures with s=1s=-1 are preferred for C-shunt flux couplers (see  [12]), we also analyze floating transmon coupler architectures that couple to qubits with opposite signs. This allows for a more direct comparison between both architectures. For the fluxonium system, we adopt the architecture described in  [14], where the transmon couplers couple to fluxonium qubits with the same sign. The local terms $\hat{H}_{\text{loc}}$ depend on the type of qubit or coupler and are discussed in the corresponding sections. They can also include disorder, which leads to many-body localization (MBL) of the system [17].
@@ II.2 MBL states in superconducting architectures
$\hat{H}_{\text{MBL}}=\sum_{\bm{n}\in\{0,1\}^{N}}w_{\bm{n}}\bigotimes_{i=1}^{N}\hat{\tau}_{i}^{n_{i}}\,$, (2)
[…]
@@
Models of superconducting qubits like Eq. (1) also fit into the above framework, with $|\tilde{\bm{n}}\rangle$ becoming the computational basis states and $\hat{\tau}_{i}$ the $Z$ operators for the dressed qubits. A difference compared with the spin-1/2 case is that the systems include higher-lying excited states for the qubits, and tunable couplers as additional sites. The computational basis states are therefore only those eigenstates that are matched with the product states in which the qubits are in one of the two lowest energy levels and the couplers are in their ground states (ignoring coupling terms between qubits or couplers). After projecting onto the $2^{N}$-dimensional computational subspace, the Hamiltonian again has the form (2), which defines the longitudinal interactions between idling qubits. The coefficients $w_{\bm{n}}$ describe how the frequency of a qubit shifts depending on the state of the other qubits and could thus be measured in experiments.
@@
In addition to these interactions, we want to investigate how well the computational basis is localized, i.e., how much the dressed basis states $|\tilde{\bm{n}}\rangle$ differ from the corresponding product states $|\bm{n}\rangle$. A possible measure for this is the inverse participatio ratio (IPR). For a state $|\psi\rangle$, it is defined as
IPR $=\sum_{\bm{n}}|\langle\psi|\bm{n}\rangle|^{4}\,$, (3)
where the sum is over all states in the undressed basis. As the interaction coefficients, the IPR can be probed in experiments as well. By applying a Rabi pulse to one qubit and testing whether a neighboring qubit has also been flipped, one could measure how well the qubit states are localized.
@@
To determine the coupling coefficients $w_{\bm{n}}$ and the IPR numerically, we make use of the fact that MBL states in one dimension admit an efficient approximation as MPS [23] and employ the DMRG-X method [16]. The difference compared to the regular DMRG method is that it targets eigenstates based on their overlap with a reference state, in this case the relevant product states $|\bm{n}\rangle$. Unless noted otherwise, the mapping between product states and eigenstates simplifies to picking the eigenstate with the largest overlap, and one DMRG-X calculation for each $n$ is sufficient. If two product states map to the same eigenstate, we do additional DMRG-X simulations with an orthogonalization against the previously calculated states, and switch to the more general mapping mentioned above that maximizes the sum of overlaps [21].
@@
The interaction coefficients are obtained from the energies $E_{\bm{n}}$ by a Walsh-Hadamard transformation [7]
$w_{\bm{n}}$ $=\frac{1}{2^{N}}\sum_{\bm{n}^{\prime}\in\{0,1\}^{N}}(-1)^{\sum_{i}n_{i}n_{i}^{\prime}}E_{\bm{n}^{\prime}}\,.$ (4)
@@
Because of the larger local Hilbert spaces, we used a single-site variant of the DMRG-X algorithm. As long as the bond dimension $\chi$ is small, the local eigenproblems in the DMRG-X can be efficiently solved by applying a standard solver to the dense-matrix representation of the projected Hamiltonians. For larger $\chi$, we switched to a matrix-free Krylov method and look for the eigenvector to the smallest eigenvalue of the $(\hat{H}-\lambda)^{2}$, where $\lambda$ is the approximate energy from a calculation with small $\chi$. We checked the convergence for each state by monitoring the variance $\langle\tilde{\bm{n}}|\hat{H}^{2}|\tilde{\bm{n}}\rangle-\langle\tilde{\bm{n}}|\hat{H}|\tilde{\bm{n}}\rangle^{2}$. In cases where the calculations did not converge to the targeted eigenstate with sufficient accuracy straightaway, usually because the DMRG simulations got stuck in a superposition of two eigenstates close in energy, we found it helpful to apply an iteration of the shift-invert DMRG method [24], implemented with the library Krylov.jl [25]. To expand the MPS bonds, we used the method described in Ref. [26].
We note that a DMRG-X approach was used recently to study lattices of superconducting qubits. In that work, a multi-target version of the algorithm was proposed to improve convergence for strongly hybridized states [27].
@@ III Transmon qubits
[…]
where $\hat{\phi}$ is the superconducting phase that obeys the commutation relation $\hat{\phi},\hat{n}]=\text{i}$ with the Cooper-pair number operator $\hat{n}$. It is convenient to also introduce the ladder operators for the quadratic part of $\hat{H}_{\text{T}}$ via
@@
$\hat{H}_{\text{T}}$ $\approx\nu\,\hat{b}^{\dagger}\hat{b}+\frac{\alpha}{2}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b}\,$, (8)
with $\nu=\sqrt{8E_{C}E_{J}}-E_{C}$ and $\alpha=-E_{C}$. In the following we truncate the Hilbert space to the lowest $6$ eigenstates and study lattices of superconducting qubits described by the Hamiltonian (1) with local terms as in Eq. (8). The charging energy is assumed to be $E_{C}=0.2\,\text{GHz}$ for all qubits, while the Josephson energy $E_{J}$ is site-dependent, introducing disorder in the qubit frequencies $\nu$. Throughout this work, we set $h=1$.
@@
$\nu_{i}$ $=\bar{\nu}+\Delta\sqrt{2}\sin\Big[\pi x_{i}\big(y_{i}+\sqrt{y_{i}^{2}+4}\big)\Big]\,.$ (9)
In an infinite ladder, $\bar{\nu}$ and $\Delta$ become the average frequency and the standard deviation, respectively. We choose $\bar{\nu}=5.8\,\text{GHz}$, $\Delta=0.33\,\text{GHz}$, and an $8$-qubit lattice with $x_{i}\in\{1,2,3,4\}$ and $y_{i}\in\{1,2\}$ (see Table 1). For the couplings between transmon qubits, we consider three different cases: the ladder (i) with only static capacitive coupling, (ii) with tunable transmons and (iii) with tunable C-shunt flux couplers.
qubit $E_{C}$ $E_{J}$ $\nu$
11 0.2 24.33 6.040
22 0.2 19.52 5.389
33 0.2 26.13 6.266
44 0.2 19.68 5.411
55 0.2 20.20 5.485
66 0.2 26.12 6.265
77 0.2 19.81 5.430
88 0.2 23.11 5.881
@@ III.1 Static capacitive coupling
Figure 2(a) displays the interaction coefficients $w_{\bm{n}}$ as a function of the capacitive-coupling strength $T$ for a system without tunable couplers ($T_{c}=0$). Naturally, the dominant $ZZ$ couplings are those between nearest neighbors, which reach values on the order of $10\,\text{kHz}$ already around $T\approx 2\,\text{MHz}$. Longer-ranged terms are orders of magnitude smaller at weak capacitive coupling, but become comparable in strength as $T$ increases, indicating that the delocalized regime is approached. In addition to stronger interactions, increasing $T$ leads to a delocalization of the dressed qubits. This is captured by the average IPR of the computational basis states, which remains close to $1$ up to around T=5T=5\,MHz before it drops off and reaches values of approximately $0.6$ for T=20T=20\,MHz [Fig.2(b)]. Overall, our results confirm previous exact-diagonalization studies of MBL in transmon arrays [7, 29, 31].
Figure 2: (a) Interaction coefficients $w_{\bm{n}}$ for transmons without couplers on a ladder. The graphical labels indicate the positions of the involved $Z$ operators (up to rotations and translations), e.g., the red curves show horizontal and vertical nearest-neighbor couplings, and the green ones the next-nearest-neighbor couplings along the diagonals. Also included are the specific longer-ranged couplings discussed in the main text (purple and orange curves). In the shaded region, the state assignment is ambiguous. The dips in some curves signal sign changes of the corresponding interaction terms. (b) Average IPR for all computational basis states and separated based on the states of the qubits $3$ and $6$.
It is expected that resonant clusters of sites form as an MBL system gets closer to the delocalization transition [32]. Here, the first sign of this appears in interaction coefficient for the next-nearest neighbor qubits 3 and 6, whose frequencies $\nu_{i}$ are almost degenerate. An indication of the strong hybridization is that the IPR is much lower for states with a single excitation between the two qubits. We also observe substantial long-ranged couplings $w_{\bm{n}}$ involving qubits 3 and 6 that correspond to the bitstrings $\bm{n}\in\{00101000,00010100,00111000,00011100\}$ in Fig. 2(a). For $T>15$ MHz, the definition of the computational basis states becomes ambiguous, as picking the eigenstate with the largest overlap is no longer an injective map. The results shown are based on the state assigment described in Sec. II. There is still an overall increase of the interaction strengths $|w_{\bm{n}}|$ with $T$, but with jumps for some coefficients due to changes in the state mapping. One could also assign the computational basis states by tracing the spectrum as a function of $T$ and identifying the anticrossings where the state labels switch [7]. Since that method is based on the energy, it might lead to a less erratic $T$-dependence of the interaction coefficients for large capacitive couplings. It is clear, however, that the system in this regime would not be suitable as a quantum-computing device and that any method of state assignment breaks down for sufficiently large $T$ when the system becomes delocalized.
@@ III.2 Tunable transmon couplers
As shown in Fig. 3, the strength of the nearest-neighbor terms is reduced by about an order of magnitude compared to the system without tunable couplers, and there is a clearer separation in energy scales to longer-ranged interactions. Specific long-ranged terms related to the qubits 3 and 6 still become significant for large enough $T$, however, indicating that these tunable couplers are not able to fully remove the harmful effect of nearly resonant qubits. At very strong couplings T=20T=20\,MHz, there are jumps in some of the interaction coefficients $w_{\bm{n}}$, likely caused by the switching of state labels at avoided level crossings [7]. Such difficulties in assigning the computational basis states are pushed to larger $T$ compared to the system without tunable couplers.
Figure 4: (a) IPR averaged over all states of the computational basis for a ladder with either tunable transmons or C-shunt flux qubits as couplers. The inset shows the average hopping amplitudes between qubits and couplers, taking the factors in Eqs. (7) and  (19) into account. (b) average mutual information between a qubit and the remaining system excluding the connected couplers.
@@
with $S(\rho)=-\text{Tr}(\rho\ln\rho)$ the von Neumann entanglement entropy and $\rho_{X}=\text{Tr}_{\bar{X}}(|\psi\rangle\langle\psi|)$ the reduced density matrix of subsystem $X$ (with complement $\bar{X}$). This quantity, averaged over all computational basis states and qubits, does not increase much until around T=10T=10\,MHz, whereas the corresponding quantity in the model without couplers already starts to rise around T=5T=5\,MHz. Tunable couplers thus appear to be effective in keeping the qubits localized in the sense that they reduce the spread of the actual (slightly hybridized) computational basis to other qubit sites.